You are given that 1+i is a root of the equation z^2+2z^2+az+b=0,where a and b are real numbers .which of the following is true?


1
Expert's answer
2021-10-25T13:36:29-0400

"z^2+2z^2+az+b=0"

"\Rightarrow 3z^2+az+b=0"

As "1+i" is a root of "3z^2+az+b=0"

So, "3(1+i)^2+a(1+i)+b=0" [Put "z=1+i" ]

"\Rightarrow" "3(1+i^2+2i)+a+ia+b=0"

"\Rightarrow 3(1-1+2i)+a+ia+b=0"

"\Rightarrow 6i+ia+a+b=0"

"\Rightarrow i(6+a)+a+b=0"

Comparing real and imaginary parts both sides, we get

"6+a=0, a+b=0
\\\Rightarrow a=-6, b=-a
\\\Rightarrow a=-6, b=6"